# Characterizing Vehicle-Induced Distributed Acoustic Sensing Signals for Accurate Urban Near-Surface Imaging

### v2

@@ 6.2 Surface wave sensitivity analysis
The depth sensitivity of surface waves is closely related to their frequency content. The Rayleigh surface-wave phase velocity sensitivity kernels, corresponding to the best model derived from heavy vehicle signals (Figure 6c), are shown in Figure 9. It can be observed that surface waves at lower frequencies (2–5 Hz) are primarily influenced by structures between 30 to 80 meters in depth. Improved estimation of phase velocities at these lower frequencies enhances sensitivity for imaging deeper structures, whereas structures above the uppermost 10 meters are more sensitive at frequencies above 10 Hz.
@@ 6.3 Insights into near-surface characterization using vehicle-induced DAS signals
Figure 9: Rayleigh surface-wave phase velocity sensitivity kernels. The color version of this figure is available only in the electronic edition.

### v3

@@ 6.2 Surface wave sensitivity analysis
The depth sensitivity of surface waves is closely related to their frequency content. The Rayleigh surface-wave phase velocity sensitivity kernels, corresponding to the best model derived from heavy vehicle signals (yellow dashed line in Figure 6c), are shown in Figure 9. It can be observed that surface waves at lower frequencies (2–5 Hz) are primarily influenced by structures between 30 to 80 meters in depth. Improved estimation of phase velocities at these lower frequencies enhances sensitivity for imaging deeper structures, whereas structures above the uppermost 10 meters are more sensitive at frequencies above 10 Hz.
@@ 6.3 Insights into near-surface characterization using vehicle-induced DAS signals
Figure 9: Rayleigh surface-wave phase velocity sensitivity kernels, corresponding to the best model derived from heavy vehicle signals (yellow dashed line in Figure 6c). The color version of this figure is available only in the electronic edition.